# Fundamentals and advances in transverse thermoelectrics

### v2

@@ VI.1 Heat flux sensing
As shown in Fig. 8(b), clear ANE signals with hysteresis behaviors were observed in the $H$ dependence of $V$ and their magnitude increases with increasing $q$ across the device. The saturated $V$ values for $\mu_{0}H>0.1~{\rm T}$ indicate the negligibly small ONE contribution, enabling the simple extraction of the ANE contribution. Figure 8(c) shows the ANE voltage, represented by $V_{\rm odd}$, as a function of $q$ [161], where $V_{\rm odd}$ at 0.2 T was extracted by the field-odd component of $V$ at $|\mu_{0}H|=0.2~{\rm T}$ and $V_{\rm odd}$ at zero field was extracted from the zero-field $V$ values for the positive-to-negative and negative-to-positive sweep of $H$. The sensitivity of the ANE-based heat flux sensor was estimated to be $V_{\rm ANE}/q\sim V_{\rm odd}/j_{\rm q}=1.12\pm 0.02\times 10^{-7}~{\rm VW}^{-1}{\rm m}^{2}$ and $0.89\pm 0.02\times 10^{-7}~{\rm VW}^{-1}{\rm m}^{2}$ in the saturation magnetization (0.2 T) and remanent magnetization (0.0 T) states, respectively. Importantly, the use of magnetic thin films with finite coercivity and remanent magnetization makes it possible to sense a heat flux using ANE even in the absence of an external magnetic field. Therefore, the design and control of magnetic anisotropy are essential for the application of ANE. Although the sensitivity of the sensor used in this demonstration is lower than that of commercial sensors based on the Seebeck effect, ANE has strong advantages in flexibility and low thermal resistance, extending applications of heat flux sensors. To further improve the performance of the ANE-based heat flux sensors, extensive efforts have been devoted to developing new materials with larger $S_{\rm ANE}$ and lower $\kappa$ as well as to optimizing the device structure. Since amorphous materials can be fabricated on any surface and have lower $\kappa$ than crystalline materials, they are promising candidates for the development of low-cost, high-performance heat flux sensors with large areas [59, 60, 61, 62, 161].

### v3

@@ VI.1 Heat flux sensing
As shown in Fig. 8(b), clear ANE signals with hysteresis behaviors were observed in the $H$ dependence of $V$ and their magnitude increases with increasing $q$ across the device. The saturated $V$ values for $\mu_{0}H>0.1~{\rm T}$ indicate the negligibly small ONE contribution, enabling the simple extraction of the ANE contribution. Figure 8(c) shows the ANE voltage, represented by $V_{\rm odd}$, as a function of $q$ [161], where $V_{\rm odd}$ at 0.2 T was extracted by the field-odd component of $V$ at $|\mu_{0}H|=0.2~{\rm T}$ and $V_{\rm odd}$ at zero field was extracted from the zero-field $V$ values for the positive-to-negative and negative-to-positive sweep of $H$. The sensitivity of the ANE-based heat flux sensor was estimated to be $V_{\rm ANE}/q\sim V_{\rm odd}/q=1.12\pm 0.02\times 10^{-7}~{\rm VW}^{-1}{\rm m}^{2}$ and $0.89\pm 0.02\times 10^{-7}~{\rm VW}^{-1}{\rm m}^{2}$ in the saturation magnetization (0.2 T) and remanent magnetization (0.0 T) states, respectively. Importantly, the use of magnetic thin films with finite coercivity and remanent magnetization makes it possible to sense a heat flux using ANE even in the absence of an external magnetic field. Therefore, the design and control of magnetic anisotropy are essential for the application of ANE. Although the sensitivity of the sensor used in this demonstration is lower than that of commercial sensors based on the Seebeck effect, ANE has strong advantages in flexibility and low thermal resistance, extending applications of heat flux sensors. To further improve the performance of the ANE-based heat flux sensors, extensive efforts have been devoted to developing new materials with larger $S_{\rm ANE}$ and lower $\kappa$ as well as to optimizing the device structure. Since amorphous materials can be fabricated on any surface and have lower $\kappa$ than crystalline materials, they are promising candidates for the development of low-cost, high-performance heat flux sensors with large areas [59, 60, 61, 62, 161].